# Psychiatry-Bench: A Multi-Task Benchmark for LLMs in Psychiatry

### v1

@@ 4.1.1 Task 1: Diagnoses
This task focuses on determining psychiatric diagnoses based on rich clinical vignettes that simulate real-world case presentations. Each entry in this section is a natural language question paragraph with a detailed patient scenario, often incorporating elements such as the patient’s psychiatric and medical history, presenting symptoms, and sometimes findings from the Mental State Examination (MSE) or contextual physical observations. The cases require models to identify either a likely diagnosis or generate a differential diagnosis, drawing on a nuanced understanding of symptom patterns, temporal progression, and psychosocial factors. For instance, cases may involve individuals with compulsive behaviors and intrusive thoughts indicative of obsessive-compulsive disorder or patients presenting with fear and avoidance consistent with specific phobias. The task challenges models to interpret and synthesize complex narratives, apply diagnostic reasoning, and align answers with formal criteria from standard psychiatric classifications such as the DSM-5 or ICD-10. As such, it plays a critical role in evaluating systems intended for clinical decision support, psychiatric triage, and medical education.
4.1.2 Task 2:Treatment
This task centers on identifying appropriate treatment strategies for various psychiatric disorders based on the clinical scenarios provided. Each case includes a natural language question focused on management, along with contextual patient information such as symptoms, diagnosis, comorbidities, or previous treatment responses. The goal is to generate or select medically accurate treatment plans that align with current psychiatric guidelines. Treatment recommendations may involve pharmacological interventions such as prescribing antidepressants, antipsychotics, mood stabilizers, or anxiolytics, as well as non-pharmacological approaches, including cognitive behavioral therapy (CBT), psychoeducation, or hospitalization in severe cases. In some instances, the model must also account for patient-specific factors, such as age, substance use, risk level, or treatment resistance. This task tests the system’s ability to apply evidence-based psychiatry, reflect clinical guidelines (e.g., NICE, APA), and personalize care based on the patient’s profile. It is particularly relevant for clinical decision support systems, therapeutic recommendation engines, and educational tools aimed at medical trainees.
4.1.3 Task 3: Treatment Follow-Up
@@
This task addresses the longitudinal management of psychiatric patients, focused on follow-up care after an initial treatment decision has been made. This subset of clinical vignettes spans multiple time points, allowing models to reason about how a patient’s mental or physical disorders evolve in response to therapy. Each entry includes an initial clinical scenario, one or more follow-up reports (e.g., new symptoms, lab results, patient response), and a question prompting clinical judgment about how care should be adjusted. These follow-up scenarios often require interpreting medication effects, identifying emerging side effects (e.g., serotonin syndrome, orthostatic hypotension), assessing treatment adherence, or recommending new interventions. In addition to psychiatric symptoms, the task incorporates age-specific considerations (e.g., geriatric psychiatry) and physical health interactions, reflecting the complexity of real-world cases. Models are expected to make informed decisions about whether to continue, adjust, or discontinue treatment. This task simulates ongoing psychiatric care, testing a model’s ability to adapt recommendations over time and handle incomplete, evolving information, making it critical for clinical decision support, chronic care planning, and training in longitudinal psychiatric case management.
4.1.4 Task 4:Classification
A multi-label dataset for mental health disorder classification is created by manually extracting case studies and their corresponding final classifications from multiple books. Each extraction undergoes a review process to ensure quality. Disorder names across different books often show inconsistencies, with similar disorders labeled under varying names. To standardize the terminology, two disorder lists are developed: a category disorders list and a specific disorders list. Both lists are extracted from the DSM-5-TR and ICD-11 manuals and are reviewed by a licensed psychiatrist to ensure accuracy. Disorder names from the books are then mapped to the standardized lists using Gemini 2.5 Pro. This mapping process is conducted twice, once using the disorder categories list and once using the specific disorders list. As a result, each case in the final dataset is assigned two sets of disorder labels: one set corresponding to the specific disorders and another set corresponding to their broader categories.
4.1.5 Task 5: Management Plan
@@
This task involves generating a structured and contextually appropriate management plan for patients presenting with psychiatric symptoms. Each entry in this portion’s detailed clinical history is often accompanied by information from a Mental State Examination (MSE) and Physical Examination, followed by a question such as "How would you manage this case?" or "What interventions should you offer?". The answers outline holistic, step-by-step care strategies, integrating diagnostic clarification, risk assessment, therapeutic planning, and follow-up. Management plans may include pharmacological treatments (e.g., SSRIs for generalized anxiety), psychological therapies (e.g., CBT for OCD), psycho-education, social interventions, and referral pathways. The task requires models to interpret multiple layers of clinical input and propose management approaches aligned with psychiatric guidelines (e.g., NICE or APA). It emphasizes the practical application of knowledge in real-world clinical decision-making and supports training for complex care coordination in psychiatric settings.
4.1.6 Task 6: Clinical Approach
This task focuses on evaluating the appropriate clinical approach to a variety of psychiatric presentations, emphasizing step-by-step assessment and decision-making processes. Each question is tied to a case scenario and asks how clinicians should approach diagnosis, risk assessment, or management in a structured and evidence-based manner. Unlike tasks that ask for final diagnoses or treatments, this task emphasizes process-oriented reasoning on how clinicians gather information, rule out critical disorders, prioritize differential diagnoses, and choose appropriate initial interventions. The dataset features cases such as evaluating major depressive disorder, schizophrenia, panic disorder, or suicidal, with questions often framed as "What is the clinical approach to…?". Answers are grounded in established psychiatric practice guidelines and typically reference core elements such as history taking, Mental State Examination (MSE), collateral information, risk stratification, and initial management priorities. This task simulates the real-world cognitive workflow of psychiatric professionals, making it essential for models designed to assist with clinical decision-making, medical education, and case-based simulation training in psychiatry.
4.1.7 Task 7: Mental QA
This task involves the definition and clarification of core psychiatric concepts, syndromes, and clinical terms. Each question in the dataset prompts a concise and medically accurate explanation of a term or concept relevant to mental health, such as "chelonian," "bizarre delusions," or the pharmacological class "selective serotonin reuptake inhibitors." These terms are foundational to understanding clinical psychiatry and are often encountered in diagnostic interviews, treatment planning, and academic contexts. The answers are short, definitional statements drawn from clinical knowledge and psychiatric literature, consistent with sources like the DSM-5, pharmacological references, or foundational psychiatry textbooks. This task is designed to enhance conceptual grounding for QA models and promote explainability, particularly in systems intended for clinical education, patient-facing applications, or knowledge base construction. It also supports retrieval-augmented models and interpretable AI by encouraging explicit reasoning about medical terminology.
4.1.8 Task 8: Sequential Question Answering
This task simulates dynamic, case-based psychiatric reasoning through a structured sequence of interrelated questions. Each scenario presents a detailed patient history, followed by a series of clinical queries that must be answered in context. The questions are designed to mirror the natural flow of psychiatric evaluation, moving from the initial differential diagnosis to evidence-based justification, etiological analysis, treatment planning, and finally prognostic estimation. For instance, a single case might begin by asking "What is the likely differential diagnosis?" followed by questions like "What information supports the diagnosis?", "What are the important etiological factors?", and "What treatment options are available?". This multi-turn format requires the model to maintain coherence, track evolving information, and apply lay clinical reasoning over time. It is beneficial for developing and evaluating QA systems that simulate real-world psychiatric interviews, medical board exam preparation, or decision-support tools in longitudinal patient care. This task emphasizes temporal context, cognitive continuity, and adaptive inference, making it a high-value benchmark for sequential understanding in medical AI.
4.1.9 Task 9: Multiple Choice Questions (MCQ)
This task includes three widely used assessment formats in psychiatric education and certification: (1) standard multiple-choice questions (MCQs), (2) extended matching items (EMIs), and (3) full-length exam simulations. Together, they are designed to evaluate knowledge recall, clinical reasoning, and decision-making in structured, test-like disorders.
  1. 1.
    Multiple Choice Questions (MCQ): The standard MCQs in our dataset are drawn from psychiatry-focused exams and cover a broad range of topics, including diagnosis, treatment, symptom classification, theoretical models, and psychopharmacology. Each question presents a clinical or conceptual prompt, followed by a list of five options, with one correct answer. These questions are designed to assess factual knowledge as well as applied understanding and clinical judgment.
  2. 2.
    Extended Matching Items (EMI): The EMI format groups several related questions under a single theme, such as diagnostic systems or therapeutic categories. Each item shares a standard answer list and must be matched with the most appropriate response. To maximize the utility of this format for QA evaluation, we split each EMI set into separate question-answer pairs while maintaining the original answer pool context. Enables fine-grained evaluation of the model’s ability to differentiate between similar options and interpret questions that depend on thematic nuance. The EMI task is particularly effective at testing classification accuracy, thematic grouping, and structured reasoning across semantically related content.
  3. 3.
    Exam Simulations: We also included complete exam sets composed of multiple MCQs delivered together, replicating the structure and pacing of real psychiatric board exams. These exams are designed to simulate actual test disorders, including variation in difficulty, topic coverage, and question style. This component allows for evaluation of performance across a complete test and supports use cases such as automated exam preparation and curriculum-aligned feedback systems.
Taken together, these three formats provide a comprehensive benchmark for evaluating model performance on structured psychiatric assessments. They also support practical applications in educational settings, such as intelligent tutoring, exam readiness, and adaptive learning.
4.2 Language Models
@@
In this section, we detail the LLMs evaluated throughout our study, covering both general-purpose and domain-specialized systems. The goal is to assess their capabilities in handling various clinically relevant tasks, such as diagnosis, treatment planning, and psychiatric reasoning. By comparing general LLMs with medical-domain-specific models, we aim to highlight the performance trade-offs and advantages of each category. This dual perspective provides a clearer understanding of how LLMs can be optimized or selected for different healthcare applications.
@@ 4.2.1 General Models
Gemini 2.5 Pro Preview 03-25 [42] Proprietary Closed-source Google’s Gemini API
Gemini 2.5 Flash Preview 04-17 (Thinking) [42] Proprietary Closed-source Google’s Gemini API
Gemini 2.5 Flash Preview 04-17 [42] Proprietary Closed-source Google’s Gemini API
Gemini 2 Flash [42] Proprietary Closed-source Google’s Gemini API
Claude Sonnet 4 [43] Proprietary Closed-source Anthropic API
DeepSeek R1 [44] 236B Open-source DeepSeek API
DeepSeek Chat [44] 67B Open-source DeepSeek API
LLaMA 3.3 70B [45] 70B Open-source Together API
QWQ 32 [46] 32B Open-source Nvidia API
QWEN 3 32b [46] 32B Open-source Deepinfra API
Table  3 presents a concise overview of the LLMs evaluated in the study. It categorizes each model by name, parameter size, source type (open or closed), and the API service provider used to access them. The models include both proprietary and open-source systems. Proprietary models, such as various versions of Google’s Gemini (e.g., Gemini 2.5 Pro, Flash Preview, and Flash), and Anthropic’s Claude Sonnet 4, are closed-source, meaning their architectures and training details are not publicly available. These models are accessed through their respective API services, such as Google’s Gemini API and the Anthropic API.
On the other hand, open-source models like DeepSeek R1 (236B), DeepSeek Chat (67B), LLaMA 3.3 70B, QWQ 32, and QWEN 3 32B offer more transparency in terms of model architecture and training datasets. These are made available through platforms such as the DeepSeek API, Together API, Nvidia API, and Deepinfra API.
The parameter sizes of the open-source models are explicitly stated, which is important for understanding their potential capabilities and computational requirements. The proprietary models do not disclose their parameter counts, reflecting the limited public insight into their underlying architectures.
Overall, the table illustrates a diverse selection of LLMs sourced from both major tech companies and open research communities, providing a broad basis for comparative experimentation and analysis.
4.2.2 Medical Models
Table 4: Comparison of Medical Language Models
Model Size (Parameters) Source API Service Provider
Llama3-OpenBioLLM-70B [47] 70B Open-source Hugging Face
Palmyra-Med-70B [48] 70B Open-source Hugging Face
Llama-MedX v3.2 [49] 70B Open-source Hugging Face
JSL-MedLlama-3-8B-v2.0 [50] 8B Open-source Hugging Face
MedGemma [51] 27B Open-source Hugging Face
II-Medical-8B [52] 8B Open-source Hugging Face
Med-PaLM (v2) [53] Proprietary Closed-source Google Cloud Vertex AI
To ensure a well-rounded and insightful evaluation of LLMs in the context of psychiatric question answering, we selected a diverse group of medical-focused models that reflect a range of architectures, parameter scales, and accessibility levels. This curation was designed to balance frontier performance with practical considerations such as openness, reproducibility, and ease of deployment.
Our model set includes high-capacity open-source systems like Llama3-OpenBioLLM-70B, Palmyra-Med-70B, and Llama-MedX v3.2, which leverage recent advances in foundational models and are optimized for biomedical and clinical tasks. We also included JSL-MedLlama-3-8B-v2.0, a smaller model aimed at providing strong performance while remaining lightweight and adaptable. These open-source models are readily accessible via platforms such as Hugging Face and are suitable for both academic research and fine-tuned clinical deployments.
@@ 4.2.2 Medical Models
4.3 Evaluation Methodology
To assess the performance and alignment of model outputs, we employ a set of evaluation metrics designed to quantify agreement, consistency, and quality across different prompts and models. These metrics allow us to systematically compare outputs, identify patterns in model behavior, and measure improvements across prompt variations. In particular, we focus on agreement counts between model pairs to capture alignment and shared reasoning under different prompting disorders.
4.3.1 Accuracy Metrics
We mainly use two accuracy metrics in this work;
  • Balanced Accuracy (BA) is used to handle imbalanced datasets by computing the average recall for each class. It is defined as follows:
    $BA=\frac{1}{2}\left(\frac{TP}{TP+FN}+\frac{TN}{TN+FP}\right)$ (1)
    BA helps in evaluating the model’s performance when the class distribution is imbalanced.
  • Accuracy is the proportion of correctly classified samples among the total number of samples:
    $Accuracy=\frac{TP+TN}{TP+TN+FP+FN}$ (2)
    Accuracy gives an overall measure of correctness, but it may not be reliable in cases of class imbalance, as it can be biased towards the majority class.
Here, $TP$ (True Positives) and $FN$ (False Negatives) are used to compute recall (sensitivity), while $TN$ (True Negatives) and $FP$ (False Positives) are used to compute specificity.

### v2

@@ 3.2.1 Diagnosis
This subset contains 467 expert-annotated question-answer pairs centered on diagnostic reasoning. The task focuses on determining psychiatric diagnoses based on rich clinical vignettes that simulate real-world case presentations. Each item includes a clinical vignette and a diagnostic query, requiring inference of the most appropriate psychiatric diagnosis. These scenarios often incorporate elements such as the patient’s psychiatric and medical history, presenting symptoms, findings from the MSE, and physical examination. The cases, representing 123 unique diagnostic questions and 295 distinct clinical histories, require models to identify either a likely diagnosis or generate a differential diagnosis, drawing on a nuanced understanding of symptom patterns and temporal progression. Examples are drawn from six authoritative sources, including 100 Cases in Psychiatry and Case Files Psychiatry. This task challenges models to interpret and synthesize complex narratives and apply diagnostic reasoning aligned with formal criteria from standard classifications like the DSM-5-TR, making it critical for evaluating systems intended for clinical decision support and medical education.
3.2.2 Treatment
This subset comprises 258 expert-annotated clinical question-answer pairs focused on treatment decision-making. The task centers on identifying appropriate treatment strategies based on clinical scenarios that include a detailed case history, a MSE, physical examination ,and a treatment-related question. The expert-curated answers reflect real-world psychiatric care and cover a wide range of disorders and treatment modalities, including pharmacological interventions (e.g., prescribing antidepressants, antipsychotics, mood stabilizers), psychotherapeutic approaches (e.g., cognitive behavioral therapy, psychoeducation, motivational interviewing), and acute care level decisions such as hospitalization when risk severity or treatment resistance requires intensive monitoring and intervention. Sourced from texts like 100 Cases in Psychiatry and Geriatric Psychiatry, this task tests a system’s ability to apply evidence-based psychiatry and personalize care according to clinical guidelines.
3.2.3 Treatment Follow-Up
@@
This dataset comprises 27 clinical question-answer pairs focused on the longitudinal management of psychiatric patients after an initial treatment decision has been made. Each scenario includes a psychiatric case history, a primary follow-up question, and often one or more follow-up reports (e.g., new symptoms, lab results), with a subset of 6 entries extending to a second-tier follow-up. These items evaluate a model’s capacity to support ongoing clinical care by interpreting medication effects, identifying emerging side effects, assessing adherence, or recommending adjustments to the treatment plan. The task simulates ongoing psychiatric care, testing a model’s ability to adapt recommendations over time and handle the incomplete, evolving information that is critical for chronic care planning.
3.2.4 Classification
The classification component was developed to evaluate the ability of LLMs to categorize psychiatric disorders based on clinical descriptions. It includes 418 psychiatric case descriptions, each annotated with both a broad diagnostic category and a specific disorder label. Clinical case samples were manually extracted from authoritative casebooks and annotated with preliminary labels. To standardize terminology, two reference lists were created from the DSM-5-TR: one for high-level diagnostic categories (e.g., "mood disorders") and another for specific disorders (e.g., "bipolar I disorder"); see Appendix 1.5A for the full lists. A mapping process was then applied to align each sample with both the validated category-level and disorder-level labels, enabling multi-label classification. Both lists were checked for clinical accuracy by a licensed psychiatrist.
3.2.5 Management Plan
@@
This dataset consists of 337 case-based question-answer pairs that test a model’s ability to generate a structured and contextually appropriate management plan. Each scenario provides a detailed clinical history, often supplemented with MSE and physical examination findings, followed by a question like "How would you manage this case?". The answers outline holistic, step-by-step care strategies that integrate diagnostic clarification, risk assessment, therapeutic planning (pharmacological and psychological), psycho-education, social interventions, and follow-up.The task emphasizes comprehensive care planning and management decision-making beyond single diagnoses or treatment choices, reflecting the complex care coordination required in real-world psychiatric settings.
3.2.6 Clinical Approach
The development of this subset began with extracting patient case scenarios from psychiatric reference books. Since these source vignettes were not accompanied by ready-made questions, it was necessary to transform them into evaluable QA pairs. For this purpose, these cases were used to generate open-ended clinical reasoning questions using the Gemini 2.5 Pro model, which was instructed to formulate questions requiring multi-step interpretive reasoning. Subsequently, each generated question underwent manual verification to ensure its clinical relevance. This subset contains 56 expertly crafted scenarios designed to evaluate a model’s ability to emulate the diagnostic thought process of a psychiatrist. Unlike tasks that ask for a final diagnosis, this component focuses on process-oriented reasoning, assessing how clinicians should gather information, rule out critical disorders, prioritize differential diagnoses, and choose initial interventions. The questions emphasize the rationale behind clinical decisions, such as identifying red flags or determining the most appropriate investigative step. This task is essential for assessing how well LLMs can replicate the nuanced cognitive workflow of trained psychiatrists. Unlike tasks that ask for final diagnoses or treatments, this task emphasizes process-oriented clinical reasoning how clinicians gather information, rule out critical disorders, prioritize differential diagnoses, and choose appropriate initial interventions.
3.2.7 Mental QA
This dataset includes 326 expert-annotated question-answer pairs that test foundational psychiatric knowledge through the definition and clarification of core concepts, syndromes, and clinical terms. Questions, sourced primarily from the Case Files and Case Vignettes series, ask models to define terms like "bizarre delusions," explain concepts such as "thought withdrawal," or describe pharmacological classes. This task targets the factual psychiatric knowledge and terminology aligned with the DSM-5-TR and formal training curricula, enhancing conceptual grounding and promoting explainability in systems intended for clinical education or knowledge base construction.
3.2.8 Sequential Question Answering
This task simulates dynamic, case-based diagnostic reasoning through a dataset of 32 clinical vignettes, each accompanied by a structured sequence of interrelated questions. These questions were designed to reflect the natural progression of a psychiatric evaluation, encompassing five core domains of clinical reasoning: (1) differential diagnosis, (2) supporting evidence, (3) etiological factors, (4) treatment options, and (5) prognosis. To extend the evaluative scope of the dataset, we introduced an additional sixth question classification based on DSM-5-TR, which is not conventionally included in standard psychiatric textbooks. This item was purposefully developed to facilitate diagnostic classification, enabling each vignette to be mapped to a unique label representing the preferred diagnosis. This multi-turn structure provides a rigorous framework for assessing a model’s capacity for longitudinal reasoning, contextual retention, and adaptive clinical judgment across sequential interactions. For example, in a vignette describing a 10-year-old boy presenting with poor eye contact, delayed speech, and repetitive behaviors, the model is guided through the reasoning sequence: identifying possible differential diagnoses (e.g., Autism Spectrum Disorder vs. Communication Disorder), outlining supporting evidence from the history, exploring potential etiological factors, proposing treatment options, and discussing prognosis. The final classification question then requires the model to determine the preferred DSM-5-TR diagnosis, which in this case would be Autism Spectrum Disorder.
3.2.9 Question Formats and Exam Simulations
This task group encompasses a variety of structured assessment formats commonly used in psychiatric board examinations and educational settings to assess knowledge recall, clinical recognition, and treatment matching. The development process for all items involved manually curating questions from reputable psychiatry resources and implementing a dedicated preprocessing step to remove all accompanying explanatory text, ensuring an unbiased evaluation format.
Standard Multiple-Choice Questions (MCQ)
This component consists of 1353 standard multiple-choice questions, each with 4-5 answer options. These items reflect the structured testing style common in psychiatric board examinations and are designed to function as clinical knowledge checks grounded in patient histories and symptom-based reasoning. For instance, a question might describe a patient with specific movements and a family history of psychiatric illness, requiring the identification of a gene location associated with a particular disease. As noted in the introductory description for this section, a dedicated preprocessing step was implemented. Many of the original source items were accompanied by detailed explanations or rationales, which, while pedagogically useful, could introduce bias in LLM evaluation. To prevent this, all explanatory text was removed, ensuring only the question stem and options were retained for an unbiased evaluation. Furthermore, all answer formats were standardized from various original formats (e.g., Roman numerals) into a consistent A/B/C/D labeling scheme, and ambiguous or duplicate questions were filtered to maintain quality.
Extended Matching Items (EMI) This task is designed to assess deeper clinical reasoning by requiring models to match multiple clinical vignettes to a single, shared list of 8-15 answer options (the "theme"). This format tests pattern recognition and fine-grained differentiation across closely related disorders or treatments. The dataset contains 277 whole EMI clusters. To support more granular evaluation and fine-tuning, the full EMI format was also disaggregated into 1037 standalone "EMI Separated" question items. Each separated item retains its original header and option list for context but functions as an independent unit, simplifying scoring and improving input granularity during training.
Exam Simulations This component includes curated exam-style datasets designed to simulate standardized testing conditions and replicate the structure of real psychiatric board exams. This set contains 600 questions split into two parts for evaluation: 300 EMI questions and 300 MCQs. These items are designed to assess factual recall and decision-making across a complete test, allowing for the evaluation of a model’s performance in a sustained, exam-like scenario.
4 Experimental and Evaluation Methodology
This section outlines the experimental framework used to evaluate the performance of LLMs across the aforementioned tasks (as defined in Section 3), which simulate/reflect real-world clinical workflows. We describe the models employed, including both general-purpose LLMs and domain-specialized medical models. The evaluation methodology (prompting, parsing, and metrics) is summarized here; details of the benchmark construction are provided in Section 3. The evaluation methodology combines key implementation details such as prompting and output parsing, automatic metrics, and LLM-as-a-judge assessments to capture performance on accuracy, reasoning depth, and clinical alignment. By structuring our experiments around tasks that simulate/reflect realistic clinical scenarios and rigorous evaluation criteria, we aim to produce insights into the capabilities and limitations of current LLMs in mental health domains.
4.1 Language Models under Evaluation
@@
In this section, we detail the LLMs evaluated throughout our study, covering both general-purpose and domain-specialized systems. The goal is to assess their capabilities in handling various clinically relevant tasks (see Section 3 for task definitions), such as diagnosis, treatment planning, and diagnostic reasoning. By comparing general LLMs with medical-domain-specific models, we aim to highlight the performance trade-offs and advantages of each category. This dual perspective provides a clearer understanding of how LLMs can be optimized or selected for different healthcare applications.
@@ 4.1.1 General Models
Model Size (Parameters) Source
API Service
Provider
Release (Year
& Quarter)
Gemini 2.5 Pro Preview 03-25 [google2024geminiapi] Proprietary Closed-source Google’s Gemini API 2025 Q1
Gemini 2.5 Flash Preview 04-17 Thinking [google2024geminiapi] Proprietary Closed-source Google’s Gemini API 2025 Q2
Gemini 2.5 Flash Preview 04-17 [google2024geminiapi] Proprietary Closed-source Google’s Gemini API 2025 Q2
Gemini 2 Flash [google2024geminiapi] Proprietary Closed-source Google’s Gemini API 2025 Q1
Claude Sonnet 4 [sonnet2024] Proprietary Closed-source Anthropic API 2025 Q2
Claude Sonnet 4.5 [sonnet2024] Proprietary Closed-source Anthropic API 2025 Q3
Claude Sonnet 4.5 Thinking [sonnet2024] Proprietary Closed-source Anthropic API 2025 Q3
GPT 5 Medium Thinking [openai2025gpt5] Proprietary Closed-source OpenAI API 2025 Q3
DeepSeek-R1 [DeepSeek2024] 671B Open-source DeepSeek API 2025 Q1
DeepSeek Chat [DeepSeek2024] 67B Open-source DeepSeek API 2024 Q4
LLaMA 3.3 70B [dubey2024llama] 70B Open-source Together API 2024 Q4
QWQ 32 [qwen2024] 32B Open-source Nvidia API 2025 Q1
Qwen 3 32B [qwen2024] 32B Open-source Deepinfra API 2025 Q2
GPT-OSS [dubey2024llama] 120B Open-source Together API 2025 Q3
Table  3 presents a concise overview of the LLMs evaluated in the study. It categorizes each model by name, parameter size, source type (open or closed), and the API service provider used to access them.The models include both proprietary and open-source systems. Proprietary models, such as various versions of Google’s Gemini (Gemini 2.5 Pro Preview, Gemini 2.5 Flash "Thinking," and Gemini 2 Flash), Anthropic’s Claude series (Claude Sonnet 4, 4.5, and 4.5 Thinking), and OpenAI’s GPT 5 Medium Thinking, are closed-source and there is no official public data on model size. As their architectures and training details are not publicly available, they are accessed through their respective API services. Models with ’Thinking’ in their name (e.g., GPT 5 Medium Thinking) may be referred to with a (T) suffix for brevity.
On the other hand, the study includes open-source models like DeepSeek-R1 (671B), DeepSeek Chat (67B), LLaMA 3.3 70B, QWQ 32 (32B), Qwen 3 32B, and GPT-OSS (120B). These models offer more transparency and are made available through platforms such as the DeepSeek API, Together API, Nvidia API, and Deepinfra API. These models have sizes ranging from 32B-671B parameters. Overall, the table illustrates a diverse selection of LLMs sourced from both major tech companies and open research communities, providing a broad basis for comparative experimentation and analysis.
4.1.2 Medical Models
Table 4: Summary of Medical Models used in our experiments
@@ 4.1.2 Medical Models
Release (Year
& Quarter)
Llama3-OpenBioLLM-70B [openbiollm2024] 70B Open-source Hugging Face 2024 Q2
Palmyra-Med-70B [palmyra_med_2024] 70B Open-source Hugging Face 2025 Q1
Llama-MedX v3.2 [llama_medx_v3_2_2025] 70B Open-source Hugging Face 2025 Q1
JSL-MedLlama-3-8B-v2.0 [jsl_medllama3_8b_v2_0_2025] 8B Open-source Hugging Face 2025 Q2
MedGemma [medgemma2025] 27B Open-source Hugging Face 2025 Q2
II-Medical-8B [2025II-Medical-8B] 8B Open-source Hugging Face 2025 Q3
Med-PaLM (v2) [medpalm_v2_2024] Proprietary Closed-source Google Cloud Vertex AI 2023 Q1
To ensure a well-rounded and insightful evaluation of LLMs in the context of psychiatry, we selected a diverse group of medical-focused models that reflect a range of architectures, parameter scales, and accessibility levels, as shown in Table 4. This curation was designed to balance frontier performance with practical considerations such as openness, reproducibility, and ease of deployment.
Our model set includes high-capacity open-source systems like Llama3-OpenBioLLM-70B, Palmyra-Med-70B, Llama-MedX v3.2, and MedGemma, which leverage recent advances in foundational models and are optimized for biomedical and clinical tasks. We also included JSL-MedLlama-3-8B-v2.0, a smaller model aimed at providing strong performance while remaining lightweight and adaptable. These open-source models are readily accessible via platforms such as Hugging Face and are suitable for both academic research and fine-tuned clinical deployments.
In addition, we explicitly report each model’s Release (Year & Quarter) to contextualize performance within the rapidly evolving landscape of medical-oriented LLMs. Because model capabilities improve significantly over short development cycles, situating each system within its temporal release window enables more meaningful comparisons. This temporal framing highlights trends such as the recent surge of high-capacity open-source biomedical models, while also clarifying how older, proprietary systems like Med-PaLM (v2) continue to serve as reference benchmarks. Including release timing therefore supports a more accurate interpretation of model maturity, recency, and relevance for clinical and educational applications.
@@
4.2 Evaluation Framework
To assess the performance and alignment of model outputs, we employ a set of evaluation metrics designed to quantify agreement, consistency, and quality across different prompts and models. These metrics allow us to systematically compare outputs, identify patterns in model behavior, and measure improvements across prompt variations. In particular, we focus on agreement counts between model pairs to capture alignment and shared reasoning under different prompting conditions.
4.2.1 Accuracy
Accuracy is the proportion of correctly classified samples among the total number of samples. It provides an overall measure of correctness; however, it may be misleading in the presence of class imbalance, as it can be biased toward the majority class.
$\text{Accuracy}=\frac{\text{Number of correct predictions}}{\text{Total number of predictions}}$ (1)